# 3D-MVP: 3D Multiview Pretraining for Robotic Manipulation

### v1

@@ Abstract
Recent works have shown that visual pretraining on egocentric datasets using masked autoencoders (MAE) can improve generalization for downstream robotics tasks [40, 29]. However, these approaches pretrain only on 2D images, while many robotics applications require 3D scene understanding. In this work, we propose 3D-MVP, a novel approach for 3D multi-view pretraining using masked autoencoders. We leverage Robotic View Transformer (RVT), which uses a multi-view transformer to understand the 3D scene and predict gripper pose actions. We split RVT’s multi-view transformer into visual encoder and action decoder, and pretrain its visual encoder using masked autoencoding on large-scale 3D datasets such as Objaverse. We evaluate 3D-MVP on a suite of virtual robot manipulation tasks and demonstrate improved performance over baselines. We also show promising results on a real robot platform with minimal finetuning. Our results suggest that 3D-aware pretraining is a promising approach to improve sample efficiency and generalization of vision-based robotic manipulation policies. We will release code and pretrained models for 3D-MVP to facilitate future research.
@@ 1 Introduction
Building learning-based manipulation systems is challenging due to the unavailability of diverse large-scale robotics data. To address this, there has been significant interest in using computer vision techniques to learn generalizable visual representations without robotics focused data, for example by self-supervised pre-training on image datasets. In particular, inspired by the success of masked language modeling in NLP, several recent works have explored masked autoencoding (MAE) for visual representation learning [19]. MAE learns to reconstruct randomly masked patches in an input image, encouraging the model to learn high-level semantic features. When applied to egocentric videos from human demonstrations, MAE has been shown to learn representations that generalize well to downstream robotics tasks such as object manipulation [40, 32, 9].
However, current MAE approaches for robotics pretrain only on 2D images, ignoring the 3D structure of the scene. Prior works in robotics have shown that methods that build an explicit 3D visual representation of the environment are more sample efficient and generalize better than those with only 2D visual representations [15, 38, 31]. Hence, in this work, we explore how we could bring the benefits of visual pretraining to robot manipulation methods that reason with explicit 3D representations.
We propose 3D-MVP, a method for 3D Multi-View Pretraining for robot manipulation. Our approach builds upon recent advances in robot manipulation. Specifically, we use the Robotic View Transformer (RVT) [15], a state-of-the-art 3D manipulation method [15]. RVT takes as input a point cloud of the scene and builds a 3D representation of the scene, using a set of fixed orthogonal “virtual” RGBD images. These RGBD images are fed through a transformer model that fuses information across views and predicts robot actions in the form of future gripper poses.
@@ 2 Related Work
Self-supervised learning. Self-supervised learning aims to learn useful representations from unlabeled data by solving pretext tasks that do not require manual annotation. Early work in this area focused on designing pretext tasks for 2D images, such as solving jagsaw puzzles [30], constrastive learning [7, 20] or joint embedding approaches [1, 2, 5, 6, 17, 43]. Most related to our work is the masked autoencoder (MAE) approach proposed by He et al. [19], which learns to reconstruct randomly masked patches in an image. MAE has been shown to learn transferable representations for object detection and segmentation tasks. Furthermore, Bachmann et al demonstrates MAE pretraining can be extended to different modalities such as semantics and depth [3]. In this work, we extend the MAE approach to multi-view 3D scenes, enabling us to learn 3D-aware representations that are useful for robotic manipulation tasks.
@@
Visual pretraining for Robotics. Visual pretraining has demonstrated impressive generalization ability on computer vision tasks. Therefore, prior works have explored whether it works for robotics tasks as well. Specifically, the robotics community has trended towards learning representations using state-of-the-art self-supervised vision algorithms on diverse interaction datasets [16, 37, 8], and finetune the network on robotics tasks [27, 29, 40, 32, 28, 9]. 3D-MVP follows the same procedure. However, existing robotics pretraining approaches typically learn a 2D visual encoder (e.g. ResNet [21] or ViT [12]), we find they are inferior than manipulation policies which do explicit 3D modeling (e.g. RVT [15], Act3D [14]). Migrating a pretrained ViT to 3D manipulation policies is nontrivial since they do not have a 2D visual encoder. In this paper, we propose 3D-MVP, which does 3D-aware pretraining on 3D manipulation policies, to fill the gap.
@@
Learning manipulation from demonstrations. Recent work has explored using transformers for multi-task manipulation policies that predict robot actions from visual and language inputs [38, 18, 26, 36, 39]. End-to-end models like RT-1 [4], GATO [34], and InstructRL [26] directly predict 6-DoF end-effector poses but require many demonstrations to learn spatial reasoning and generalize to new scenes. To better handle 3D scenes, PerAct [38] and C2F-ARM [24] voxelize the workspace and detect the 3D voxel containing the next end-effector pose. However, precise pose prediction requires high-resolution voxels which are computationally expensive. Recently, RVT [15] proposes a multi-view transformer that attends over point cloud features from multiple camera views to predict actions. This avoids explicit voxelization and enables faster training and inference than PerAct. Act3D [14] represents the scene as a continuous 3D feature field and samples points to featurize with attention, allowing adaptive resolution. GNFactor [42] jointly optimizes a generalizable neural field for reconstruction and a Perceiver for decision-making, using a shared 3D voxel representation enriched with semantics from a vision-language model. In contrast, our proposed 3D-MVP learns 3D scene representations through masked autoencoding pretraining on a large dataset of 3D object models. This pretraining enables 3D-MVP to build a rich understanding of 3D geometry and semantics prior to finetuning on downstream manipulation tasks. Compared to RVT and Act3D which train from scratch on target tasks, 3D-MVP’s pretraining leads to improved performance, sample efficiency and generalization. Unlike GNFactor which relies on a pretrained vision-language model to inject semantics, 3D-MVP directly learns 3D semantic features from object models.
@@ 3 Approach
It is a state-of-the-art object manipulation method [15]. It creates an explicit 3D representation of the scene by using orthogonal virtual views of a scene. Please refer to Goyal et al. [15] for a full explanation. Here, we provide a brief overview and define the notation.
RVT takes a point cloud of the robot workspace as input (Fig. 1). RVT is agnostic to the poses of the RGBD cameras used to construct the input point cloud. For example, it can be obtained from a combination of third-person cameras around the workspace, head cameras, or wrist cameras. RVT then renders this point cloud using a set of five “virtual" cameras placed at orthogonal locations around the robot. The virtual cameras are placed at the top, left, right, front, and back of the robot workspace with respect to the robot. Each virtual image has 10 channels: RGB (3 channels), Depth (1 channel), 3D point coordinate in world frame (3 channels), and 3D point coordinate in camera sensor frame (3 channels). We denote the virtual images captured from different virtual camera poses $\{p_{1},\ldots,p_{5}\}$ as $\{I_{1},\ldots,I_{5}\}$.
@@
RVT is trained end-to-end from scratch on sampled trajectories from simulator or real robots. While RVT has shown state-of-the-art results on 3D manipulation, it does not generalize to novel objects and scenes. In the next section, we describe our novel approach 3D-MVP, and how we modify and pretrain the RVT encoder using 3D-MVP.
@@
Implementation details. We implement the pretraining using the PyTorch library and train it on eight NVIDIA V100 GPUs. We use the Objaverse dataset for pretraining [11], which contain a total of 800K+ 3D objects with realistic textures and materials. We sample 200K high-quality 3D models as the training set, and 1000 for validation purpose. We do not construct a large-scale validation set, since the validation is qualitative.
We use a patch size of 10x10 to tokenize images. For the encoder $\mathcal{E}$, we use a multiview transformer with 8 layers, 8 attention heads and a hidden dimension of 1024. For the decoder, we use a multiview transformer with 2 layers, with the same number of attention heads and hidden dimension. We train the model for 15 epochs using the AdamW optimizer with a learning rate of 0.0001 and a weight decay of 0.01. We use a batch size of 3 and a masking probability of 0.75.
@@ 3.3 Finetuning on Downstream Manipulation Tasks
Implementation details. For finetuning on manipulation demonstrations, we follow the standard practice [15, 31]. We use 8 NVIDIA V100 (32GB) for finetuning and a single V100 for evaluation. The learning rate is 1e-4 with warmup. We use Lamb [41] as the optimizer and the batch size is 3.
Figure 2: MAE Reconstruction results on Objaverse. We find the our pretrained multi-view transformer could generalize to unseen object instances and reconstruct multi-view images from their masked versions.
@@ 4 Experiments
In this section, we evaluate the effectiveness of 3D-MVP for robotic manipulation tasks, and aim to answer the following questions: (1) Can masked autoencoding effectively learn useful representations from multi-view 3D data? (2) Does 3D-aware pretraining improve manipulation performance compared to training from scratch or 2D pretraining? (3) Does 3D-aware pretraining improve robustness to environmental variations encountered in manipulation tasks? (4) How do various design choices while pretraining affect the downstream manipulation performance?
To answer these questions, we evaluate 3D-MVP on two benchmarks: RLBench [23] for general manipulation performance and COLOSSEUM [31] for systematic evaluation of robustness to environmental perturbations.
@@ 4.1 Validating 3D Masked Autoencoding
4.2 Results on RLBench
We then evaluate whether our proposed pretraining improves manipulation performance. The experiments are conducted on a simulated platform called RLBench [23]. Setup. RLBench [23] is a popular simulation benchmark for learning manipulation policies. Each task requires the robot to perform a specific action, such as picking up an object, opening a drawer, or stacking blocks. We follow the simulation setup of PerAct [38] and RVT [15] and use CopppelaSim [35] to simulate 18 RLBench tasks. A Franka Panda robot with a parallel gripper is controlled to complete the tasks. The 18 RLBench tasks are the same as PerAct and RVT. The visual observations are captured from four noiseless RGB-D cameras positioned at the front, left shoulder, right shoulder, and wrist with a resolution of 128×128.
Baselines. We compare 3D-MVP with the following baselines on RLBench: (1) Image-BC [25] is an image-to-action behavior cloning approach which takes the visual observation and predict the corresponding action. We compare with two variants which use CNN and ViT as the visual encoders, and call them Image-BC (CNN) and Image-BC (ViT), respectively; (2) C2F-ARM-BC [24] is another behavior cloning approach which converts RGB-D obversations to multi-resolution voxels and predicts the next key-frame action. (3) PerAct [38]: a multi-task a Perceiver transformer for robotic manipulation. The inputs are point clouds with color features and PerAct uses a Perceiver IO network to compress them to a fixed dimension [22]. (4) RVT [15]: The same Robotic View Transformer architecture as 3D-MVP but trained from scratch on the downstream tasks. We do not compare with 2D pretraining methods since they do not work well on RLBench [31].
@@ 4.2 Results on RLBench
Results. We show quantitative results on Table 1. For the average success rate, 3D-MVP outperforms existing state-of-the-art methods, PerAct [38] and RVT [15], by a large margin. It demonstrates the effectiveness of bringing visual pretraining for manipulation policy which has explicit 3D modeling. We find the improvement of 3D-MVP mainly comes from tasks which has medium difficulty, such as “insert peg”, “put in cupboard”, and “stack blocks”. If the task is too hard and PerAct/RVT is not able to solve any of them, the pretraining does not help. If the task is too easy and PerAct/RVT has already reached $>$90 success rate, the pretraining has limited space of improvement.
@@ 4.3 Results on COLOSSEUM
After validating our proposed pretraining is helpful for robotic manipulation, we further evaluate its generalization ability and robustness to environmental variations. Therefore, we evaluate 3D-MVP and the baselines on the COLOSSEUM axes of variation.
@@
Benchmark. COLOSSEUM [31] is a benchmark for evaluating generalization for robotic manipulation. It contains 20 different tasks such as hockey, empty dishwasher. For each task, it generates 12 environmental perturbations, including changes in color, texture, size of objects and backgrounds, and lightnings, distractors and camera poses. The objects which can be changed include Manipulation object (MO), Receiver Object (RO) and the table. Therefore, it is well-suited for evaluating the generalization ability of manipulation approaches with pretraining.
Simulation setup. For simulation, we follow the original COLOSSEUM setup. We use CopppelaSim [35] to simulate all tasks. In training, we do not add any environmental perturbations and generate 100 demonstrations for each task. During test time, we generate 12 environmental perturbations for each task. For each environmental perturbation, we generate 25 demonstrations. For each demonstration, we repeat the generation 20 times if it fails. However, we still experienced a few rendering errors. For example, we find it’s hard to find the right perturbation parameters for the task of “empty dishwasher”. Therefore, we only report results on settings we are able to render. We report baselines on exactly the same setting, to make sure the comparison is fair.
Metrics. We also report the task completion success rate on COLOSSEUM. Instead of reporting the average success rate for each individual tasks, we report the average success rate for each pertubation of the environment, as it will highlight how each method is robust to different pertubations.
Baselines. We compare 3D-MVP with state-of-the-art baselines reported on COLOSSEUM, which includes RVT and two 2D pretraining approaches.
$\bullet$ MVP [40, 32]: A 2D pretraining approach using MAE reconstruction losses. It is pretrained on a collection of interaction datasets, such as Ego4D [16], EpicKitchen [8], and 100DOH [37]. The pretrained encoder is then finetuned and evaluated on COLOSSEUM.
$\bullet$ R3M [29] A 2D pretraining approach using a combination of reconstruction and contrastive losses. It is pretrained on Ego4D videos and languages [16]. The pretrained encoder is then finetuned and evaluated on COLOSSEUM.
$\bullet$ RVT [15]: Trained on COLOSSEUM tasks from scratch.
Results. We show results in Figure 3. First, our method outperforms existing 2D pretraining (MVP [40], and R3M) significantly. It indicates existing 2D pretrainig methods are not ready for complicated robotic manipulation. Compared with RVT which is trained from scratch, our method is more robust to most perturbations. It is especially robust to the change of texture and size of Receiver Object (RO), size of the manipulation object (MO), Light color and Table color. We believe it is because the pretraining stage enables our approach to see diverse 3D objects.
4.4 Ablation Studies

### v2

# 3D-MVP: 3D Multiview Pretraining for Manipulation

@@ Abstract
Recent works have shown that visual pretraining on egocentric datasets using masked autoencoders (MAE) can improve generalization for downstream robotics tasks. However, these approaches pretrain only on 2D images, while many robotics applications require 3D scene understanding. In this work, we propose 3D-MVP, a novel approach for 3D Multi-View Pretraining using masked autoencoders. We leverage Robotic View Transformer (RVT), which uses a multi-view transformer to understand the 3D scene and predict gripper pose actions. We split RVT’s multi-view transformer into visual encoder and action decoder, and pretrain its visual encoder using masked autoencoding on large-scale 3D datasets such as Objaverse. We evaluate 3D-MVP on a suite of virtual robot manipulation tasks and demonstrate improved performance over baselines. Our results suggest that 3D-aware pretraining is a promising approach to improve generalization of vision-based robotic manipulation policies.
**footnotetext: The work was done during internship at NVIDIA.
@@ 1 Introduction
Building learning-based manipulation systems is challenging due to the unavailability of diverse large-scale robotics data. To address this, there has been significant interest in using computer vision techniques to learn generalizable visual representations without robotics focused data, for example by self-supervised pre-training on image datasets. In particular, inspired by the success of masked language modeling in Natural Language Processing (NLP), several recent works have explored masked autoencoding (MAE) for visual representation learning [22, 3, 13]. MAE learns to reconstruct randomly masked patches in an input image, encouraging the model to learn high-level semantic features. When applied to egocentric videos from human demonstrations, MAE has been shown to learn representations that generalize well to downstream robotics tasks such as object manipulation [42, 33, 9].
Figure 1: Comparison of 2D vs. 3D pretraining for manipulation. (Left): In 2D pretraining [42], the model is trained to do MAE reconstruction from a single image from Interaction videos [17, 39]. The encoder is then used for downstream manipulation tasks. However, the input can only be a single 2D image due to the pretraining. (Right): We propose 3D Multi-View Pretraining (3D-MVP), which uses multiple orthogonal RGB-D views of a 3D model. The model is tasked with reconstructing the masked views by leveraging information across different perspectives, enabling it to learn more robust 3D spatial representations. This multi-view approach improves downstream performance in robot manipulation tasks by capturing richer scene understanding compared to 2D-only pretraining. And it is compatible with state-of-the-art 3D manipulation method such as RVT [16] which takes 3D inputs.
However, current MAE approaches for robotics pretrain only on 2D images, ignoring the 3D structure of the scene. Prior works in robotics have shown that methods that build an explicit 3D visual representation of the environment are more sample efficient and generalize better than those with only 2D visual representations [16, 40, 32]. Hence, in this work, we explore how we can bring the benefits of visual pretraining to robot manipulation methods that reason with explicit 3D representations.
We propose 3D-MVP, a method for 3D Multi-View Pretraining for robot manipulation. Our approach builds upon recent advances in robot manipulation. Specifically, we use the Robotic View Transformer (RVT), a state-of-the-art 3D manipulation method [16]. RVT takes as input a point cloud of the scene and builds a 3D representation of the scene, using a set of fixed orthogonal “virtual” RGBD images. These RGBD images are fed through a transformer model that fuses information across views and predicts robot actions in the form of future gripper poses.
@@ 2 Related Work
Self-supervised learning. Self-supervised learning aims to learn useful representations from unlabeled data by solving pretext tasks that do not require manual annotation. Early work in this area focused on designing pretext tasks for 2D images, such as solving jagsaw puzzles [31], constrastive learning [7, 21] or joint embedding approaches [1, 2, 5, 6, 18, 45]. Most related to our work is the masked autoencoder (MAE) approach proposed by He et al. [22], which learns to reconstruct randomly masked patches in an image. MAE has been shown to learn transferable representations for object detection and segmentation tasks. Furthermore, Bachmann et al demonstrates MAE pretraining can be extended to different modalities such as semantics and depth [3]. In this work, we extend the MAE approach to multi-view 3D scenes, enabling us to learn 3D-aware representations that are useful for robotic manipulation tasks. Unlike MultiMAE which learns semantics and depth through direct supervision, 3D-MVP aims to learn a 3D-aware representation from multi view images.
@@
Visual pretraining for Robotics. Visual pretraining has demonstrated impressive generalization ability on computer vision tasks. Therefore, prior works have explored whether it works for robotics tasks as well. Specifically, the robotics community has trended towards learning representations using state-of-the-art self-supervised vision algorithms on diverse interaction datasets [17, 39, 8], and finetune the network on robotics tasks [28, 30, 42, 33, 29, 9, 37]. 3D-MVP follows the same procedure. However, existing robotics pretraining approaches typically learn a 2D visual encoder (e.g. ResNet [20] or ViT [12]), we find they are inferior than manipulation policies which do explicit 3D modeling (e.g. RVT [16], Act3D [15]). Migrating a pretrained ViT to 3D manipulation policies is nontrivial since they do not have a 2D visual encoder. In this paper, we propose 3D-MVP, which does 3D-aware pretraining on 3D manipulation policies, to fill the gap.
@@
Learning manipulation from demonstrations. Recent work has explored using transformers for multi-task manipulation policies that predict robot actions from visual and language inputs [40, 19, 27, 38, 41]. End-to-end models like RT-1 [4], GATO [35], and InstructRL [27] directly predict 6-DoF end-effector poses but require many demonstrations to learn spatial reasoning and generalize to new scenes. To better handle 3D scenes, PerAct [40] and C2F-ARM [25] voxelize the workspace and detect the 3D voxel containing the next end-effector pose. However, precise pose prediction requires high-resolution voxels which are computationally expensive. Recently, RVT [16] proposes a multi-view transformer that attends over point cloud features from multiple camera views to predict actions. This avoids explicit voxelization and enables faster training and inference than PerAct. Act3D [15] represents the scene as a continuous 3D feature field and samples points to featurize with attention, allowing adaptive resolution. GNFactor [44] jointly optimizes a generalizable neural field for reconstruction and a Perceiver for decision-making. In contrast, our proposed 3D-MVP learns 3D scene representations through masked autoencoding pretraining on a large dataset of 3D object models. This pretraining enables 3D-MVP to build a rich understanding of 3D geometry and semantics prior to finetuning on downstream manipulation tasks. Compared to RVT and Act3D which train from scratch on target tasks, 3D-MVP’s pretraining leads to improved performance, sample efficiency and generalization. Unlike GNFactor which relies on a pretrained VLM to inject semantics, 3D-MVP directly learns 3D semantic features from object models.
@@ 3 Approach
RVT is a state-of-the-art object manipulation method [16]. It creates an explicit 3D representation of the scene by using orthogonal virtual views of a scene. Please refer to Goyal et al. [16] for a full explanation. Here, we provide a brief overview and define the notation.
RVT takes a point cloud of the robot workspace as input (Fig. 2). RVT is agnostic to the poses of the RGB-D cameras used to construct the input point cloud. For example, it can be obtained from a combination of third-person cameras around the workspace, head cameras, or wrist cameras. RVT then renders this point cloud using a set of five “virtual" cameras placed at orthogonal locations around the robot. The virtual cameras are placed at the top, left, right, front, and back of the robot workspace with respect to the robot. Each virtual image has 10 channels: RGB (3 channels), Depth (1 channel), 3D point coordinate in world frame (3 channels), and 3D point coordinate in camera sensor frame (3 channels). We denote the virtual images captured from different virtual camera poses $\{p_{1},\ldots,p_{5}\}$ as $\{I_{1},\ldots,I_{5}\}$.
@@
RVT is trained end-to-end from scratch on sampled trajectories from simulator or real robots. While RVT has shown state-of-the-art results on 3D manipulation, it does not generalize to novel objects and scenes, since we train RVT from scratch and it overfits on training demonstrations. In the next section, we describe our novel approach 3D-MVP, and how we modify and pretrain the RVT encoder using 3D-MVP, to learn a generalizable representation.
@@
Implementation details. We implement the pretraining using the PyTorch library and train it on 8 NVIDIA V100s. We use the Objaverse dataset for pretraining [10], which contain a total of 800K+ 3D objects with realistic textures and materials. We sample 200K high-quality 3D models. and 1000 for validation purpose. We do not construct a large-scale validation set, since the validation is only qualitative. The evaluation of pretrained representation should be mainly done on downstream manipulation tasks.
We use a patch size of 10x10 to tokenize images. For the encoder $\mathcal{E}$, we use a multiview transformer with 8 layers, 8 attention heads and a hidden dimension of 1024. For the decoder, we use a multiview transformer with 2 layers and 8 attention heads. We train the model for 15 epochs using the AdamW optimizer with a learning rate of 0.0001 and a weight decay of 0.01. We use a batch size of 3 and a masking probability of 0.75.
Figure 3: MAE Reconstruction results on Objaverse. Our pretrained multi-view transformer generalizes to unseen object instances and reconstructs multi-view images from their masked versions.
@@ 3.3 Finetuning on Downstream Manipulation Tasks
Implementation details. For finetuning on manipulation demonstrations, we follow the standard practice [16, 32] to ensure fair comparison with baselines with 2D pretraining and without pretraining. We use 8 NVIDIA V100 (32GB) for finetuning and a single V100 for evaluation. The learning rate is 1e-4 with 2000 warmup steps and cosine learning rate decay until 1e-6. We use Lamb [43] as the optimizer and the batch size is 3. We train the model for 15 epochs on both RLBench and COLOSSEUM training set.
@@ 4 Experiments
In this section, we evaluate the effectiveness of 3D-MVP for robotic manipulation tasks, and aim to answer the following questions: (1) Does 3D-aware pretraining improve manipulation performance compared to training from scratch or 2D pretraining? (2) Does 3D-aware pretraining improve robustness to environmental variations encountered in manipulation tasks? (3) How do various design choices while pretraining affect the downstream manipulation performance? To answer these questions, we evaluate 3D-MVP on two benchmarks: RLBench [24] for general manipulation performance and COLOSSEUM [32] for systematic evaluation of robustness to environmental perturbations.
@@ 4.1 Validating 3D Masked Autoencoding
4.2 Results on RLBench
We then evaluate whether our proposed pretraining improves manipulation performance. The experiments are conducted RLBench, a simulated platform [24].
Setup. RLBench [24] is a popular simulation benchmark for learning manipulation policies. Each task requires the robot to perform a specific action, such as picking up an object, opening a drawer, or stacking blocks. We follow the simulation setup of PerAct [40] and RVT [16] and use CopppelaSim [36] to simulate 18 RLBench tasks. A Franka Panda robot with a parallel gripper is controlled to complete the tasks. The 18 RLBench tasks are the same as PerAct and RVT. The visual observations are captured from four noiseless RGB-D cameras positioned at the front, left shoulder, right shoulder, and wrist with a resolution of 128×128.
Baselines. We compare 3D-MVP with the following baselines on RLBench:
(1) Image-BC [26] is an image-to-action behavior cloning approach which takes the visual observation and predict the corresponding action. We compare with two variants which use CNN and ViT as the visual encoders, and call them Image-BC (CNN) and Image-BC (ViT), respectively;
(2) C2F-ARM-BC [25] is another behavior cloning approach which converts RGB-D obversations to multi-resolution voxels and predicts the next key-frame action.
(3) PerAct [40]: a multi-task a Perceiver transformer for robotic manipulation. The inputs are point clouds with color features and PerAct uses a Perceiver IO network to compress them to a fixed dimension [23].
(4) RVT [16]: The same Robotic View Transformer architecture as 3D-MVP but trained from scratch on the downstream tasks. We do not compare with 2D pretraining methods since they do not work well on RLBench [32].
@@ 4.2 Results on RLBench
Results. We show quantitative results on Table 1. For the average success rate, 3D-MVP outperforms existing state-of-the-art methods, PerAct [40] and RVT [16]. It demonstrates the effectiveness of bringing visual pretraining for manipulation policy which has explicit 3D modeling. We find the improvement of 3D-MVP mainly comes from tasks which has medium difficulty, such as “insert peg”, “put in cupboard”, and “stack blocks”. If the task is too hard and PerAct/RVT is not able to solve any of them, the pretraining does not help. If the task is too easy and PerAct/RVT has already reached more than 90% success rate, the pretraining has limited space of improvement.
@@ 4.3 Results on COLOSSEUM
After validating our proposed pretraining is helpful for manipulation, we further evaluate its generalization ability and robustness to environmental variations.
Benchmark. COLOSSEUM [32] is a benchmark for evaluating generalization for robotic manipulation. It contains 20 different tasks such as hockey, empty dishwasher. For each task, it has 12 environmental perturbations, including changes in color, texture, size of objects and backgrounds, and lightnings. The objects which can be changed include Manipulation object (MO), Receiver Object (RO) and the table. Results on COLOSSEUM have also shown strong corelation with real robot experiments. Therefore, it is well-suited and comprehensive for evaluating the generalization ability of manipulation approaches with pretraining on COLOSSEUM benchmark [32].
@@
Simulation setup. For simulation, we follow the original COLOSSEUM setup. We use CoppelaSim [36] to simulate all tasks. In training, we do not add any environmental perturbations and generate 100 demonstrations for each task. During test time, we generate 12 environmental perturbations for each task. For each environmental perturbation, we generate 25 demonstrations. For each demonstration, we repeat the generation 20 times if it fails. In some cases, it is not possible to generate a plausible perturbation for some scenarios. For example, we find it’s hard to find the right perturbation parameters for the task of “empty dishwasher”. Therefore, we only report results on settings we are able to generate. We report baselines on exactly the same setting to make sure the comparison is fair.
Metrics. We also report the task completion success rate on COLOSSEUM. Instead of averaging for each task, we report the average success rate for each environmental perturbation as it will highlight how each method is robust to different perturbations.
Baselines. We compare 3D-MVP with state-of-the-art baselines reported on COLOSSEUM, which includes RVT and two 2D pretraining approaches:
(1) MVP [42, 33]: a 2D pretraining approach using MAE reconstruction losses. It is pretrained on a collection of interaction datasets, such as Ego4D [17], EpicKitchen [8], and 100DOH [39]. The pretrained encoder is then finetuned and evaluated on COLOSSEUM.
(2) R3M [30]: a 2D pretraining approach using a combination of reconstruction and contrastive losses. It is pretrained on Ego4D [17]. The pretrained encoder is then finetuned and evaluated on COLOSSEUM.
(3) RVT [16]: Trained on COLOSSEUM from scratch.
Results. We show results in Figure 4. First, our method outperforms existing 2D pretraining (MVP [42], and R3M [30]) significantly. It indicates existing 2D pretraining methods are not ready for complicated robotic manipulation. Compared with RVT which is trained from scratch, our method is more robust to most perturbations. It is especially robust to the change of texture and size of Receiver Object (RO), size of the manipulation object (MO), Light color and Table color. We believe it is because the pretraining stage enables our approach to see diverse 3D objects.
4.4 Ablation Studies